3x^2-7x+2=0
A) Solve the Equation
B) Verify that the sum of solutions is -b/a
C) Verify the product of solutions is c/a

Answer by ewatrrr(24785)

3x^2-7x+2=0 Factoring
(3x - 1)(x - 2)= 0
x is 1/3, 2
1/3 + 2 = 7/3 = -b/a = -(-7)/3
(1/3)2 = 2/3 = c/a = 2/3